OpenStudy (anonymous):

solve: 5a-3/2-3a-9/8=6

OpenStudy (dumbcow):

multiply everything by 8 40a -12 -24a -9 = 48 add like terms 16a -21 = 48 16a = 69 a = 69/16

OpenStudy (anonymous):

nicely done dumb cow, i like your way better than mine! also i like your name (better than mine). :p

OpenStudy (dumbcow):

haha thanks

OpenStudy (anonymous):

thats doesnt really make sure to me? its suppose to come out as an even number

OpenStudy (dumbcow):

oh ok i think i misunderstood the question, use parenthesis it helps \[\frac{5a-3}{2} - \frac{3a-9}{8} = 6\] multiply equation by 8 leaving \[4(5a-3) - (3a-9) = 48\] distribute and add like terms \[20a-3a -12 +9 = 48\] \[17a = 51\] \[a=3\]